# Citation Benchmark

### v1

@@ Abstract
Literature search questions, such as “where can I find research on the evaluation of consistency in generated summaries?” pose significant challenges for modern search engines and retrieval systems. These questions often require a deep understanding of research concepts and the ability to reason over entire articles. In this work, we introduce LitSearch, a retrieval benchmark comprising 597 realistic literature search queries about recent ML and NLP papers. LitSearch is constructed using a combination of (1) questions generated by GPT-4 based on paragraphs containing inline citations from research papers and (2) questions about recently published papers, manually written by their authors. All LitSearch questions were manually examined or edited by experts to ensure high quality. We extensively benchmark state-of-the-art retrieval models and also evaluate two LLM-based reranking pipelines. We find a significant performance gap between BM25 and state-of-the-art dense retrievers, with a 24.8% difference in absolute recall@5. The LLM-based reranking strategies further improve the best-performing dense retriever by 4.4%. Additionally, commercial search engines and research tools like Google Search perform poorly on LitSearch, lagging behind the best dense retriever by 32 points. Taken together, these results show that LitSearch is an informative new testbed for retrieval systems while catering to a real-world use case. [^1] Our dataset and code are available at https://github.com/princeton-nlp/LitSearch.
@@ 1 Introduction
Finding literature via a specific search query—for example, to collect related work, to check if a method has been proposed before, or to recall a previously-seen paper—is a critical task for researchers. Developing systems that recommend citations pertinent to such inquiries has the promise to enhance researchers’ productivity and expedite scientific discovery Färber and Jatowt (2020). However, this task is inherently challenging as it often requires deep domain expertise and reasoning through lengthy papers.
@@
Figure 1: Examples of inline-citation questions and author-written questions from LitSearch. The questions are often challenging and require deep understanding of target papers to answer correctly.
@@
Figure 2: The pipeline to generate inline-citation questions. We first sample a citation mention and prompt GPT-4 to write a question. Then we filter the questions by word overlap (with the target paper title) and manual inspection. The manual inspection annotates the specificity and the quality of the questions (rubrics in Table 1).
@@
Prior to this study, the task of citation recommendation is often formalized by using inline citation mentions from existing papers as queries, and the cited papers as targets He et al. (2010); Gu et al. (2022). For instance, given the citation mention “RoBERTa and T5 are based on recent advances in masked language modeling [citation],” the text surrounding the citation mention is used as a retrieval query, and the cited paper is the target literature. However, directly using inline citations often leads to queries that are noisy, broad (e.g., “Large Language Models [citation]”), or context-dependent (e.g., “We follow the hyperparameters of [citation]”).
@@
In this work, we propose a new litearture retrieval benchmark called LitSearch. As illustrated in Figure 1, a literature search question seeks papers that meet specific criteria, closely reflecting actual research workflows. LitSearch consists of two subsets: (1) For inline-citation questions, we sample citation mentions from a collection of scientific papers and use GPT-4 (OpenAI, 2023) to rewrite them into literature search questions (Figure 2). We retain questions with a low word overlap with the title of the target papers, and perform manual examination to ensure high quality. (2) For author-written questions, we invited authors of ACL 2023 and ICLR 2024 papers to write literature search questions for their own papers. This subset is also manually examined and filtered to remove any inaccurate or easy questions. LitSearch contains 597 questions in total, each paired with one or more scientific papers as the ground truth.
@@
LitSearch has several unique characteristics: (1) To our best knowledge, LitSearch is the first dataset featuring realistic literature search questions, providing a new testbed for citation recommendation and retrieval systems. (2) LitSearch is challenging, requiring deep understanding and reasoning over entire articles. The average document length (6,041/134 words for full text/titles and abstracts) is significantly longer than most existing retrieval benchmarks (e.g., 56 for Nguyen et al., 2017). (3) LitSearch is of high quality, with all questions manually examined by the authors of this work.
@@
We conduct extensive experiments on both state-of-the-art retrieval models and reranking with large language models (LLMs). On LitSearch, the best dense retrieval model, GritLM (Muennighoff et al., 2024), achieves an average recall@5 of 74.8%, beating BM25 (Robertson et al., 2009) by 24.8%. The recall@5 of GritLM is further improved by 4.4% with GPT-4 reranking. On the other hand, commercial search engines and research tools like Google Search perform poorly on this task, only achieving an average recall@5 of 42.8% at most. Futhermore, compared to existing retrieval datasets from BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2022), LitSearch effectively reflects the performance differences among various embedding models, making it an informative testbed for evaluating state-of-the-art retrieval systems.
@@ 3.2 Baselines
GPT-4 reranking (w/ BM25) 54.9 60.0 67.5 77.1 76.8 82.9 59.9 68.0
GPT-4 one-hop (w/ BM25) 62.0 64.1 71.6 74.3 73.5 77.7 64.8 68.6
GPT-4 reranking (w/ GritLM) 74.7 73.2 79.9 77.1 85.8 92.4 75.3 79.2
GPT-4 one-hop (w/ GritLM) 72.9 70.3 78.4 74.3 84.4 87.2 73.2 77.0
Table 3: Main experiment results of LitSearch. Here we only use the titles and abstracts of papers for retrieval and reranking. We report recall@20 (R@20) for broad questions and recall@5,20 (R@5, R@20) for specific questions. “Broad” and “specific” correspond to the annotations during our manual filtering stage (defined in Table 1).
Figure 3: We demonstrate detailed retrieval results using BM25, E5 and GritLM up to $k=50$. Additionally, we show the effect of applying GPT-4 reranking over GritLM retrieval results.
@@
In addition to vanilla retrieval, we also use strong LLMs (GPT-4555We use gpt-4-1106-preview in our experiments. in our case) to rerank the top retrieved results from the above retrievers. We use two strategies:
Vanilla reranking. We include the top-$n$ retrieved papers (titles and abstracts) in the context and prompt GPT-4 to rerank these based on the question (see our prompt in Table 13). This is similar to prior works (Sun et al., 2023b; Ma et al., 2023). We use $n=100$, resulting in an average context length of 13,844 words.
@@
We extract the top $m$ retrieved papers and construct a new candidate list by adding papers cited by each of these seed retrieved papers. We concatenate papers in the following order, skipping duplicates: [rank-1 paper $p_{1}$, papers cited by $p_{1}$, rank-2 paper $p_{2}$, papers cited by $p_{2}$, …, $p_{m}$, papers cited by $p_{m}$]. To avoid very long contexts, we truncate this list after the first $n$ papers and use the same prompt for GPT-4 based reranking as above. In our experiments, we use $m=50$ and $n=200$ resulting in average length of 27,544 words.
@@ 3.3 Results
We also report the performance improvement brought by the reranking methods on the weakest (BM25) and strongest (GritLM) retrievers in Table 3. We observe that both vanilla and one-hop reranking improve over the base retrieval performance. For example, on the specific subset of inline questions, the vanilla GPT-4 reranking improves the recall@5 of BM25 and GritLM by $21.5\%$ and $5.5\%$ respectively. Interestingly, the improvements from one-hop reranking are generally lower than vanilla reranking across all subsets, when using GritLM as the base retriever; this shows that our benchmark cannot be easily “gamed” by mimicking the data collection pipeline or exploiting similar citation mentions to the question from other papers.
@@ 4.3 Performance of Search Engines
In practice, researchers use search engines like Google Search, Google Scholar, or Elicit [^7] https://elicit.com/ to search for relevant papers for their scientific queries. We conduct a human study to understand how these search engines perform on LitSearch: We randomly sample 80 questions (all specific; 40 inline-citation and 40 author-written) from our dataset. We manually input these questions into the above search engines and report recall@5.888For Google Search, we consider the top-5 academic papers in the search results and ignore other webpage results. We only consider the academic papers on the first page of search results, even if this number is lower than 5. We note that this is not an apples-to-apples comparison against the retrieval models in earlier sections due to the discrepancy in the retrieval corpus.
@@ 5 Related Work
The community has propose a number of citation recommendation datasets Färber and Jatowt (2020), including global citation recommendation datasets (directly using a paper as the query and papers it cites as target papers; Cohan et al., 2020; Bhagavatula et al., 2018), and local citation recommendation datasets (using inline citation mentions as queries;  He et al., 2010; Medić and Snajder, 2020; Jeong et al., 2020; Gu et al., 2022). There are also language models and retrieval models specifically trained for scientific document understanding and retrieval tasks, such as SciBERT Beltagy et al. (2019) and SPECTER Cohan et al. (2020). Compared to existing citation recommendation datasets, LitSearch is comprised of manually annotated, natural language literature search questions, providing a more realistic and challenging evaluation for citation recommendation systems.
@@
There have been numerous datasets evaluating retrieval systems from Wikipedia (Kwiatkowski et al., 2019; Lee et al., 2019), web queries (Nguyen et al., 2017), biomedical questions (Voorhees and Tice, 2000), and more. Recently, there have been several benchmarks combining multiple datasets and evaluating retrieval or embedding models across different domains and different use cases, such as KILT (Petroni et al., 2021), BEIR (Thakur et al., 2021), and MTEB (Muennighoff et al., 2022). LitSearch offers a unique perspective by exploring the novel literature search question type, effectively complementing the existing benchmarks.
@@ 6 Conclusion
In this paper, we propose LitSearch, a new retrieval benchmark comprised of 597 manually-curated literature search questions. LitSearch includes a inline-citation question set and an author-written question set, both undergoing manual inspections from the authors of LitSearch. We conduct extensive experiments with BM25, state-of-the-art embedding models, and LLM reranking. Our experiments demonstrate the superior performance of state-of-the-art instruction-finetuned embedding models, with additional improvement via GPT-4-based reranking. We also verify that commercial search engines like Google struggle on LitSearch questions. The comparison with existing retrieval benchmarks shows that LitSearch better highlights the different performance of retrieval systems.
@@
The research artifact of this paper, LitSearch, is manually inspected and is ensured to have no unsafe or inappropriate content. However, the process to generate the dataset may introduce certain biases: for example, the inline-citation questions contain more target papers that have high citations due to the sampling; the author-written questions only cover ACL 2023 and ICLR 2024 papers.
@@ Appendix E Prompts and Additional Statistics
Table 12 shows the prompt we use for generating inline-citation questions via GPT-4. Table 13 shows the reranking prompt for GPT-4. Table 14 shows a more detailed statistics about LitSearch.
@@
Table 13: The prompt used for reranking retrieved documents using GPT4 (adapted from Sun et al., 2023a).

### v2

@@ Abstract
Literature search questions, such as “Where can I find research on the evaluation of consistency in generated summaries?” pose significant challenges for modern search engines and retrieval systems. These questions often require a deep understanding of research concepts and the ability to reason across entire articles. In this work, we introduce LitSearch, a retrieval benchmark comprising 597 realistic literature search queries about recent ML and NLP papers. LitSearch is constructed using a combination of (1) questions generated by GPT-4 based on paragraphs containing inline citations from research papers and (2) questions manually written by authors about their recently published papers. All LitSearch questions were manually examined or edited by experts to ensure high quality. We extensively benchmark state-of-the-art retrieval models and also evaluate two LLM-based reranking pipelines. We find a significant performance gap between BM25 and state-of-the-art dense retrievers, with a 24.8% absolute difference in recall@5. The LLM-based reranking strategies further improve the best-performing dense retriever by 4.4%. Additionally, commercial search engines and research tools like Google Search perform poorly on LitSearch, lagging behind the best dense retriever by up to 32 recall points. Taken together, these results show that LitSearch is an informative new testbed for retrieval systems while catering to a real-world use case. [^1] Our dataset and code are available at https://github.com/princeton-nlp/LitSearch.
@@ 1 Introduction
Finding literature via a specific search query—for example, collecting related work, checking if a method has been proposed before, or recalling a previously seen paper—is a critical task for researchers. Developing systems that recommend citations pertinent to such inquiries holds the potential to enhance researchers’ productivity and expedite scientific discovery Färber and Jatowt (2020). However, this task is inherently challenging as it often requires deep domain expertise and reasoning through lengthy papers.
@@
Figure 1: Examples of inline-citation and author-written questions from LitSearch. These questions are often challenging and require a deep understanding of the target papers to answer correctly.
@@
Figure 2: The pipeline for generating inline-citation questions. We first sample a citation mention and prompt GPT-4 to generate a question. Next, we filter questions based on word overlap with the target paper title and perform manual inspections to annotate their specificity and quality (see rubrics in Table 1).
@@
Prior to this study, the task of citation recommendation was often formalized by using inline citation mentions from existing papers as queries, and the cited papers as targets He et al. (2010); Gu et al. (2022). For instance, given the citation mention “RoBERTa and T5 are based on recent advances in masked language modeling [citation],” the text surrounding the citation mention is used as a retrieval query, and the cited paper is the target literature. However, directly using inline citations often leads to queries that are noisy, overly broad (e.g., “Large Language Models [citation]”), or highly context-dependent (e.g., “We follow the hyperparameters of [citation]”).
@@
In this work, we propose a new literature retrieval benchmark called LitSearch. As illustrated in Figure 1, a literature search question seeks papers that meet specific criteria, closely reflecting actual research workflows. LitSearch consists of two subsets: (1) For inline-citation questions, we sample citation mentions from a collection of scientific papers and use GPT-4 (OpenAI, 2023) to rewrite them into literature search questions (Figure 2). We retain questions with a low word overlap with the title of the target papers, and perform manual examination to ensure high quality. (2) For author-written questions, we invited authors of ACL 2023 and ICLR 2024 papers to write literature search questions for their own papers. This subset is also manually examined and filtered to remove any inaccurate or easy questions. LitSearch contains 597 questions in total, each paired with one or more scientific papers as the ground truth.
@@
LitSearch has several unique characteristics: (1) To the best of our knowledge, LitSearch is the first dataset featuring realistic literature search questions, providing a new testbed for citation recommendation and retrieval systems. (2) LitSearch is challenging, requiring deep understanding and reasoning over entire articles. The average document length (6,041/134 words for full texts/titles and abstracts) is significantly longer than that of most existing retrieval benchmarks (e.g., 56 for (Nguyen et al., 2017)). (3) LitSearch is of high quality, with all questions manually examined by the authors.
@@
We conduct extensive experiments on both state-of-the-art retrieval models and reranking with large language models (LLMs). On LitSearch, the best dense retrieval model, GritLM (Muennighoff et al., 2024), achieves an average recall@5 of 74.8%, outperforming BM25 (Robertson et al., 2009) by 24.8%. The recall@5 of GritLM is further improved by 4.4% with GPT-4o reranking. On the other hand, commercial search engines and research tools like Google Search perform poorly on this task, only achieving an average recall@5 of 42.8% at most. Furthermore, compared to existing retrieval datasets from BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2022), LitSearch effectively reflects the performance differences among various embedding models, making it an informative testbed for evaluating state-of-the-art retrieval systems.
@@ 3.2 Baselines
GPT-4o reranking (w/ BM25) 54.9 60.0 67.5 77.1 76.8 82.9 59.9 68.0
GPT-4o one-hop (w/ BM25) 62.0 64.1 71.6 74.3 73.5 77.7 64.8 68.6
GPT-4o reranking (w/ GritLM) 74.7 73.2 79.9 77.1 85.8 92.4 75.3 79.2
GPT-4o one-hop (w/ GritLM) 72.9 70.3 78.4 74.3 84.4 87.2 73.2 77.0
Table 3: Main experimental results of LitSearch. Here we only use the titles and abstracts of papers for retrieval and reranking. We report recall@20 (R@20) for broad questions and recall@5 and @20 (R@5, R@20) for specific questions. “Broad” and “specific” correspond to the annotations during our manual filtering stage (defined in Table 1).
Figure 3: We demonstrate detailed retrieval results using BM25, E5 and GritLM up to $k=50$. Additionally, we show the effect of applying GPT-4o reranking over GritLM retrieval results.
@@
In addition to vanilla retrieval, we also use strong LLMs (GPT-4o [^5] We use gpt-4o-2024-05-13 in our experiments. in our case) to rerank the top retrieved results from the above retrievers. We use two strategies:
Vanilla reranking. We include the top-$n$ retrieved papers (titles and abstracts) in the context and prompt GPT-4o to rerank these based on the question (see our prompt in Table 13). This is similar to prior works (Sun et al., 2023b; Ma et al., 2023). We use $n=100$, resulting in an average context length of 13,844 words.
@@
We extract the top $m$ retrieved papers and construct a new candidate list by adding papers cited by each of these seed retrieved papers. We concatenate papers in the following order, skipping duplicates: [rank-1 paper $p_{1}$, papers cited by $p_{1}$, rank-2 paper $p_{2}$, papers cited by $p_{2}$, …, $p_{m}$, papers cited by $p_{m}$]. To avoid very long contexts, we truncate this list after the first $n$ papers and use the same prompt for GPT-4o based reranking as above. In our experiments, we use $m=50$ and $n=200$ resulting in average length of 27,544 words.
@@ 3.3 Results
We also report the performance improvement brought by the reranking methods on the weakest (BM25) and strongest (GritLM) retrievers in Table 3. We observe that both vanilla and one-hop reranking improve over the base retrieval performance. For example, on the specific subset of inline questions, the vanilla GPT-4o reranking improves the recall@5 of BM25 and GritLM by $21.5\%$ and $5.5\%$ respectively. Interestingly, the improvements from one-hop reranking are generally lower than vanilla reranking across all subsets, when using GritLM as the base retriever; this shows that our benchmark cannot be easily “gamed” by mimicking the data collection pipeline or exploiting similar citation mentions to the question from other papers.
@@ 4.3 Performance of Search Engines
In practice, researchers use search engines like Google Search, Google Scholar, or Elicit [^7] https://elicit.com/ to search for relevant papers for their scientific queries. We conduct a human study to understand how these search engines perform on LitSearch: We randomly sample 80 questions (all specific; 40 inline-citation and 40 author-written) from our dataset. We manually input [^8] We use incognito mode browser sessions, but do not specially prevent IP-address-based location tracking. these questions into the above search engines and report recall@5.999For Google Search, we consider the top-5 academic papers in the search results and ignore other webpage results. We only consider the academic papers on the first page of search results, even if this number is lower than 5. We note that this is not an apples-to-apples comparison against the retrieval models in earlier sections due to the discrepancy in the retrieval corpus.
@@ 5 Related Work
The community has proposed a number of citation recommendation datasets Färber and Jatowt (2020), including global citation recommendation datasets (directly using a paper as the query and papers it cites as target papers; Cohan et al., 2020; Bhagavatula et al., 2018), and local citation recommendation datasets (using inline citation mentions as queries;  He et al., 2010; Medić and Snajder, 2020; Jeong et al., 2020; Gu et al., 2022). There are also language models and retrieval models specifically trained for scientific document understanding and retrieval tasks, such as SciBERT Beltagy et al. (2019) and SPECTER Cohan et al. (2020). Compared to existing citation recommendation datasets, LitSearch is comprised of manually annotated, natural language literature search questions, providing a more realistic and challenging evaluation for citation recommendation systems.
@@
There have been numerous datasets evaluating retrieval systems from Wikipedia (Kwiatkowski et al., 2019; Lee et al., 2019), web queries (Nguyen et al., 2017), biomedical questions (Voorhees and Tice, 2000), and more. Recently, there have been several benchmarks combining multiple datasets and evaluating retrieval or embedding models across different domains and different use cases, such as KILT (Petroni et al., 2021), BEIR (Thakur et al., 2021), and MTEB (Muennighoff et al., 2022). LitSearch offers a unique perspective by exploring the novel literature search question type, effectively complementing the existing benchmarks. Contemporary with our work, CiteME (Press et al., 2024) introduces a benchmark for identifying references based on claims made in a paper’s inline text (as opposed to research questions in our case). Their focus differs from ours in that it is aimed more at LLM-based agents rather than retrieval systems.
@@ 6 Conclusion
In this paper, we propose LitSearch, a new retrieval benchmark comprising 597 manually-curated literature search questions. LitSearch includes an inline-citation question set and an author-written question set, both undergoing manual inspection from the authors of LitSearch. We conduct extensive experiments with BM25, state-of-the-art embedding models, and LLM reranking. Our experiments demonstrate the superior performance of state-of-the-art instruction-finetuned embedding models, with additional improvement via GPT-4o-based reranking. We also verify that commercial search engines like Google struggle with LitSearch questions. The comparison with existing retrieval benchmarks shows that LitSearch better differentiates the performance of retrieval systems.
@@
The research artifact of this paper, LitSearch, is manually inspected and has been ensured to have no unsafe or inappropriate content. However, the process to generate the dataset may introduce certain biases: for example, the inline-citation questions contain more target papers that have high citations due to the sampling; the author-written questions only cover ACL 2023 and ICLR 2024 papers.
@@ Appendix E Prompts and Additional Statistics
Table 12 shows the prompt we use for generating inline-citation questions via GPT-4. Table 13 shows the reranking prompt for GPT-4o. Table 14 shows a more detailed statistics about LitSearch.
@@
Table 13: The prompt used for reranking retrieved documents using GPT-4o (adapted from Sun et al., 2023a).